# Structure-Preserving Model Order Reduction for Nonlinear DAE Models of Power Networks

### v1

@@ A-C Power System Algebraic Equations and Loads Dynamics
Appendix B Proposed MOR algorithms
The proposed SP-POD and SP-BPOD-based MOR algorithms are detailed in this section. Further explanations about creating the snapshot matrices ($\bm{X}_{d}$, $\bm{X}_{a}$, and $\bm{X}_{f}$) and covariance matrices ($\bm{G}_{c_{11}}$ and $\bm{G}_{o_{11}}$) are given in Sec. VI.
3 Create snapshot matrices $\bm{X}_{d}$ (5a), $\bm{X}_{a}$ (5b), and $\bm{X}_{f}$ (10)
4 if $n<<t$ then
5      Perform SVD as:
7else
8      Perform eigenvalue decomposition:
10       Then design $\bm{W}_{d}$ and $\bm{W}_{a}$ as folows:
12Construct $\bm{W}=\mathrm{blkdiag}(\bm{W}_{d},\;\bm{W}_{a})$ as in (6)
14 Construct $\bm{T}=\mathrm{blkdiag}(\bm{T}_{d},\;\bm{T}_{a})$ as in (7)
15 Design $\bm{W}_{R}=\bm{T}\bm{W}$ and $\bm{W}_{L}=\bm{W}_{R}^{-1}$
16 SVD $\bm{X}_{f}$ (10) as: $\bm{X}_{f}=\bm{W}_{f}\bm{\Sigma}_{f}\bm{\Lambda}_{f}$
17 Start greedy algorithm to design $\bm{P}_{M}$
18 Construct rank-$r$ approximating basis
$\bm{W}_{fr}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{r}}\end{bmatrix}$
19 Choose the first index: $[\rho$, $i_{1}]$ $=\mathrm{max}(\bm{w}_{f_{1}})$
20 Construct first measurement matrix
$\bm{P}_{M_{1}}=\bm{e}_{i_{1}},\,\bm{W}_{fr}=[\bm{w}_{f_{1}}]$
21 for $j=2:r$ do
22       calculate $\bm{c}$ using, $\bm{P}_{M}^{\top}\bm{W}_{fr}\bm{c}=\bm{P}_{M}\bm{w}_{f_{j}}$
23       compute residual, $\bm{d}=\bm{w}_{f_{j}}-\bm{W}_{fr}\bm{c}$
24       update $\bm{P}_{M}$ and $\bm{W}_{fr}$ as follows:
25       $[\rho,i_{j}]=\mathrm{max}(\bm{d})$
26       $\bm{W}_{fr}=[\bm{W}_{fr},\bm{w}_{f_{j}}],\,\bm{P}_{M}=[\bm{P}_{M},\bm{e}_{i_{j}}]$
Algorithm 1 SP-POD for power system NDAE models
3 Compute matrices $\bm{G}_{c_{11}}$ and $\bm{G}_{o_{11}}$ using (15) and (16)
4 Compute $\bm{T}_{1}$ (20), $\bm{T}_{2}$ (23), $\bm{T}_{3}$ (24), and $\bm{T}_{4}$ (25)
5 Construct $\bm{W}_{d}=\bm{T}_{1}\bm{T}_{2}\bm{T}_{3}\bm{T}_{4}$ as in (26)
7 Select $\bm{W}_{a}=\bm{W}_{gc}$
8 Compute $\bm{W}=\mathrm{blkdiag}(\bm{W}_{d},\;\bm{W}_{a})$ as in (6)
10 Design matrix $\bm{T}=\mathrm{blkdiag}(\bm{T}_{d},\;\bm{T}_{a})$ as in (7)
11 Design $\bm{W}_{L}=\bm{W}\bm{T}$ and $\bm{W}_{R}=\bm{W}_{L}^{-1}$
13 Start greedy algorithm to design $\bm{P}_{M}$
14 Construct rank-$r$ approximating basis
$\bm{W}_{fr}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{r}}\end{bmatrix}$
15 Choose the first index: $[\rho$, $i_{1}]$ $=\mathrm{max}(\bm{w}_{f_{1}})$
16 Initialize measurement matrix
$\bm{P}_{M_{1}}=\bm{e}_{i_{1}},\,\bm{W}_{fr}=[\bm{w}_{f_{1}}]$
17 for $j=2:r$ do
18       calculate $\bm{c}$ using, $\bm{P}_{M}^{\top}\bm{W}_{fr}\bm{c}=\bm{P}_{M}\bm{w}_{f_{j}}$
19       compute residual, $\bm{d}=\bm{w}_{f_{j}}-\bm{W}_{fr}\bm{c}$
20       update $\bm{P}_{M}$ and $\bm{W}_{fr}$ as follows:
21       $[\rho,i_{j}]=\mathrm{max}(\bm{d})$
22       $\bm{W}_{fr}=[\bm{W}_{fr},\bm{w}_{f_{j}}],\,\bm{P}_{M}=[\bm{P}_{M},\bm{e}_{i_{j}}]$
Algorithm 2 SP-BPOD for power system NDAE models
Figure 6: Diagram of the IEEE 39-bus system with solar power plant at Buses $36$ and a motor load at Bus $14$ while the symbol G represents buses with synchronous generators.

### v2

@@ A-C Power System Algebraic Equations and Loads Dynamics
[Uncaptioned image] Muhammad Nadeem was born in Nilore, Islamabad, Pakistan. He received the B.E. and M.S. degrees in Electrical Engineering Power from Air University and the National University of Science and Technology (NUST) Islamabad Pakistan in 2017 and 2020. He is currently a graduate research assistant pursuing a Ph.D. degree in civil and environmental engineering at Vanderbilt University, Nashville, Tennessee, USA. His research interests include control theory, reinforcement learning, system identification, and state estimation for renewables integrated power systems.
[Uncaptioned image] Ahmad F. Taha is an associate professor with the Department of Civil and Environmental Engineering at Vanderbilt University in Nashville, Tennessee. He has a secondary appointment in Electrical and Computer Engineering. He received his B.E. and Ph.D. degrees in Electrical and Computer Engineering from the American University of Beirut, Lebanon in 2011 and Purdue University, West Lafayette, Indiana in 2015. Prior to joining Vanderbilt. Taha was an assistant professor with the ECE department at the University of Texas, San Antonio (UTSA). Dr. Taha is interested in understanding how complex cyber‐physical, urban infrastructure operate, behave, and occasionally misbehave. His research focus includes optimization, control, monitoring, and security of infrastructure with applications to power, water, and transportation systems. Dr. Taha is an associate editor of IEEE Transactions on Control of Network Systems.